OpenStudy (anonymous):

Solve for x: X-5/4 = 3. 17 7 12 10

thomaster (thomaster):

\(\Large\dfrac{x-5}{4}=3\) ?

OpenStudy (anonymous):



thomaster (thomaster):

Oke :) start with multiplying both sides by 4

OpenStudy (anonymous):

as in 4*3 And 4*4

thomaster (thomaster):

\(\Large\dfrac{x-5}{4}=3\) both sides multiplied by 4: the /4 gets canceled out and 3*4 = 12 so: \(\Large x-5=12\)

thomaster (thomaster):

I believe you know how to do the last step :)

OpenStudy (anonymous):

no refresh my memory

thomaster (thomaster):

how do we get rid of the -5 on the left? we need to get it to the right

OpenStudy (anonymous):